Activity: Amortization Payment

4.   UCPB is offering mortgages at 9% interest. What

monthly payments would be required to amortize a loan of P2,000,000 for 15

years?


5. A debt of P40,000 is to be amortized by equal

payments at the end of every quarter for 1 ½ years. If the interest charged is

12% compounded quarterly, find the outstanding principal after each payment is

made. For this item, construct an amortization table.


1
Expert's answer
2021-05-24T18:55:58-0400

4) r=9%,n=15,A=2000000


Using the formula-


"A=S\times \dfrac{r}{(1+r)^n-1}\\[9pt]\Rightarrow 2000000=S\times \dfrac{0.09}{(1+0.09)^{15}-1}\\[9pt]\Rightarrow 2000000=S\times \dfrac{0.09}{2.642}\\[9pt]\Rightarrow S=2000000\times \dfrac{2.642}{0.09}=58711111.11"


Hence Monthly payment is 58711111.11



5) Calculate quarterly payment for debt as shown below:

Quarterly payment "А=S×( \dfrac{r}{
(1+r) ^n−1})"

Quarterly rate 3.00% 12%/4

No of payments 6 1.5*4

Quarterly payment

"=\\dfrac{ 4000}{(1+0\n.03)^6-1)}\t\t\n\\\\[9pt]\n\t\t\t\t\n\n =\\dfrac{ 40000}{5.4172}"


=7383.92

Prepare amortization table as shown below: